Ammonia (NH3, MM 17.031 g/mol) and hypobromite (OBr−) react according to the following chemical reaction.

2NH3+3OBr−⟶N2+3Br−+3H2O

Suppose 0.809 g of ammonia reacts with 37.2 mL of a hypobromite solution. Calculate the concentration of the hypobromite solution.


1
Expert's answer
2022-04-08T07:16:04-0400

Moles ammonia = 0.809 / 17.031 = 0.0475 (mol)

The mole ratio is 2 NH3 : 3 OBr-.

So, moles OBr- = 3/2 * 0.0475 = 0.071 (mol)

Concentration = 0.071 / 0.0372 = 1.92 M